OpenStudy (starlord1):

Use a transformation to solve the equation. t – (–16.5) = –13.4 A. t = –29.9 B. t = –3.1 C. t = 3.1 D. t = 29.9

OpenStudy (anonymous):

i think A but i am not sure i am not good at math

OpenStudy (anonymous):

Your best guess is mine lol

OpenStudy (starlord1):

that is correct

OpenStudy (anonymous):

it is?

OpenStudy (anonymous):

well that's good

OpenStudy (mathmate):

To know if it is correct, substitute t=-29.9 in t-(-16.5) = -29.9+16.5=-13.4 so good!